OpenStudy (lxelle):

The function f is such that f(x) = 2 sin^2x - 3 cos^2x for 0 <= x <= p. (i) Express f(x) in the form a + b cos^2x, stating the values of a and b. (ii) State the greatest and least values of f(x). (iii) Solve the equation f(x) + 1 = 0.

OpenStudy (lxelle):

I have found i) a=2, b=-5.

OpenStudy (dumbcow):

for part 2) set derivative equal to 0 f(x) = 2 - 5cos^2 f'(x) = -10cos x (-sinx) = 10 sin cos = 0 split up factors sin x = 0 ----> x = 0,pi cos x =0 ---> x = pi/2 f(0) = 2-5 = -3 f(pi/2) = 2 - 0 = 2 max = 2, min = -3

OpenStudy (dumbcow):

part 3) f(x) +1 = 0 3 - 5cos^2 = 0 cos^2 x = 3/5 cos x = +- sqrt(3/5) take inverse cos you should get 4 solutions

OpenStudy (lxelle):

Is there other simpler way to get ii) answer. I dont get why must it be differentiated.

OpenStudy (dumbcow):

haha yes i guess there is because its a trig function cos^2 is bounded from 0 to 1 just plug in these values for cos^2 to get max,min values

OpenStudy (lxelle):

what's the graph for cos^2x

OpenStudy (lxelle):

thankssss.